kekeman:

Find the total surface area of a cylinder with a height of 7 cm and radius of 3 cm. Leave your answer in terms of π. 140π cm2 91π cm2 60π cm2 51π cm2

snowflake0531:

\[Surface~area~of~cylinder:~~ 2\pi r^2 + 2\pi r(h)\]

snowflake0531:

So you have \[2\pi(9)+2\pi(3)(7) = 18\pi + 42\pi\] and then you can get your answer

kekeman:

60π cm2

snowflake0531:

yes